# Test-Time Scaling of Diffusion Models via Noise Trajectory Search

### v1

@@ Abstract
The iterative and stochastic nature of diffusion models enables test-time scaling, whereby spending additional compute during denoising generates higher-fidelity samples. Increasing the number of denoising steps is the primary scaling axis, but this yields quickly diminishing returns. Instead optimizing the noise trajectory—the sequence of injected noise vectors—is promising, as the specific noise realizations critically affect sample quality; but this is challenging due to a high-dimensional search space, complex noise-outcome interactions, and costly trajectory evaluations. We address this by first casting diffusion as a Markov Decision Process (MDP) with a terminal reward, showing tree-search methods such as Monte Carlo tree search (MCTS) to be meaningful but impractical. To balance performance and efficiency, we then resort to a relaxation of MDP, where we view denoising as a sequence of independent contextual bandits. This allows us to introduce an $\epsilon$-greedy search algorithm that globally explores at extreme timesteps and locally exploits during the intermediate steps where de-mixing occurs. Experiments on EDM and Stable Diffusion reveal state-of-the-art scores for class-conditioned/text-to-image generation, exceeding baselines by up to $164\%$ and matching/exceeding MCTS performance. To our knowledge, this is the first practical method for test-time noise trajectory optimization of arbitrary (non-differentiable) rewards.
@@ 1 Introduction
Figure 1: (Left) Implicit denoising tree traversed by search algorithms. (Right) Visualization of local search in noise space to maximize reward at a single timestep ${t}$.
@@ 2 Related Work
While test-time scaling has been predominantly explored for language models (Yao et al., 2023; Muennighoff et al., 2025), recent work has begun to investigate test-time scaling behavior for diffusion. One means of investing more compute in diffusion model sampling is increasing the number of denoising steps; this generally leads to better generations, but with diminishing benefits, due to accumulation of approximation and discretization errors (Xu et al., 2023). In this work, we fix the denoising steps and explore scaling along an orthogonal axis: compute devoted to searching for more favorable noise candidates during the sampling process.
@@
Zhang et al. (2025) propose T-SCEND, which sequentially performs best-of-$N$ random search and MCTS as denoising proceeds; however this paper does not truly implement standard MCTS as (1) at timestep $t$ they estimate reward by computing the predicted clean sample $\hat{\mathbf{x}}_{0}^{(t)}$ via Tweedie’s formula (Efron, 2011) (which estimates the posterior mean $\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t},\mathbf{c}]$ from a noisy sample $\mathbf{x}_{t}$, see Appendix B) rather than doing a complete rollout and then backpropagating reward on the true clean sample, making their method still a greedy search; and (2) their method performs best-of-$N$ search for the majority of denoising steps, only doing their “MCTS” for a small number ($\sim$3333-$5$) of denoising steps near the very end of the sampling process. This version of MCTS can no longer be considered a true approximation of exhaustive tree search.
@@ 3.2 From Markov Decision Processes to Contextual Bandits
A continuous contextual bandit problem can be viewed as a specialized single-state, one-step MDP with a continuous action space. It is defined by a tuple $(\Phi,\mathcal{A},\tilde{R})$, where $\Phi$ is the space of contexts, $\mathcal{A}$ is the continuous action space (shared with the definition of MDP), and $\tilde{R}$ is the reward function specific to the bandit setting. In each round $i$, the agent observes a context $\phi_{i}\in\Phi$ (analogous to an initial state $\mathbf{s}_{0}$ from $\rho_{0}$ in an MDP, but where contexts for different rounds are typically drawn independently). The agent then selects an action $\mathbf{a}_{i}\in\mathcal{A}$ based on its policy $\pi(\mathbf{a}_{i}\mid\phi_{i})$ and immediately receives a reward $\tilde{r}_{i}=\tilde{R}(\phi_{i},\mathbf{a}_{i})$.
@@ 4.2 Monte-Carlo Tree Search (MCTS)
MCTS approximates exhaustive search over the $N^{T}$ possible noise trajectories, with the approximation error going to 00 as the number of simulations $S\to\infty$. Full MCTS though faces two key drawbacks: (1) it is extremely computationally expensive, with its number of function evaluations (NFEs)—i.e. number of forward passes through $D_{\theta}$—scaling quadratically in the number of denoising steps $T$ (see Table 3 for each sampling method’s NFE formula); and (2) MCTS by definition is restricted to a fixed denoising tree, i.e. at each timestep $t$ we cannot explore “noise space” beyond the $N$ initially determined candidates. This is a critical issue, as the initial $N$ candidates are sampled uniformly from our prior (i.e., $\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$)) as opposed to any higher-reward regions of noise space that might exist. In other words, we have no notion of importance sampling; we are allocating the entire computation cost of running MCTS to samples that in general, likely yield lower-than-optimal rewards (Chatterjee and Diaconis, 2017).
@@ 4.3 Noise Search via Contextual Bandits
Recall that the SDE diffusion formulation requires the injected noise samples at each timestep to be independent (and Normally distributed). Thus, if we assume independence of actions across timesteps, and ignore the impact of the action at timestep $t$ on the state at $t-1$, we can reformulate denoising as a sequence of $T$ independent continuous contextual bandit problems; at each timestep $t$, we have the bandit characterized by:
@@ 4.4 ${\epsilon}$-greedy
Inspired by the semi-uniform strategies originally proposed to approximately solve bandit problems (Slivkins, 2024), including the original $\epsilon$-greedy bandit algorithm (Sutton and Barto, 1998), we introduce a novel sampling algorithm that is an $\epsilon$-contaminated version of zero-order search (henceforth simply “$\epsilon$-greedy” (Alg. 1), where at each timestep $t$: (1) We start with a random Gaussian noise $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ as our pivot. (2) We then find $N$ candidates $\{\mathbf{z}_{t}^{(i)}\}_{i=1}^{N}$, via the following procedure: with probability $\epsilon$, $\mathbf{z}_{t}^{(i)}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$; else, we draw a random $\mathbf{z}_{t}^{(i)}$ from the pivot’s neighborhood ($\lambda\sqrt{2d}$-radius ball around $\mathbf{p}$). (3) We compute the clean sample predicted $\hat{\mathbf{x}}_{0}^{(i)}$ using Tweedie’s formula using each candidate $\mathbf{z}_{t}^{(i)}$. (4) Finally, we score each clean sample, set the pivot to the noise candidate leading to the clean sample with the highest scoring, and repeat steps 1-3 $K-1$ times, finally setting $\mathbf{z}_{t}$ to the last value of $\mathbf{p}$.
@@
1:Discretization timesteps $T$, context (e.g. class label) $\mathbf{c}$, learned denoising network $D_{\theta}$, max. step size scaling factor $\lambda$, number of local search iterations $K$, branching factor (number of noise candidates) $N$, sampling step function $f$, mixture proportion $\epsilon$, reward function $r$, initial noise sample $\mathbf{x}_{T}$
2:for $t=T$ to $1$ do Sample $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ // pivot
@@
yields remarkable performance improvements, which we discuss in section 5.
@@ 5.1.3 Results
Table 1: EDM results by sampling method. Each value in columns 2-4 is obtained by generating $36$ images given random ImageNet class labels. The column header denotes the reward used both to score the final images and during sampling if applicable. We set $\lambda=0.15$ and $\epsilon=0.4$. Note that we measure distance as $\|\mathbf{z}_{t}^{(i)}-\mathbf{p}\|_{F}$, hence let $\lambda$ a scaling factor applied to $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{N}(0,1)^{d}}\left[\|\mathbf{x}-\mathbf{y}\|_{F}\right]\approx\sqrt{2d}$ (where $\|\cdot\|_{F}$ is the Frobenius norm. Re. notation, $N$ is branching factor, $B$ is beam width, $S$ is number of MCTS simulations, and $K$ is the number of local search iterations. Generating each sample took $<$$1$ second for naive sampling (the lowest-compute method) and $<$1 minute for MCTS (the highest-compute method) on a single A100 (40GB). We provide error bars for each reward and sampling method, computed as the standard deviation of scores for a given prompt over $20$ generations with variability from randomness of the noise draws.
@@
Naive Sampling $0.4965$±$\pm 0.01$ $0.3563$±$\pm 0.07$ $0.3778$±$\pm 0.04$ $18$
Best of $4$ Sampling $0.5767$±$\pm 0.01$ $0.4220$±$\pm 0.02$ $0.5461$±$\pm 0.00$ $72$
Beam Search ($N=4,B=2$) $0.6334$±$\pm 0.02$ $0.4679$±$\pm 0.05$ $0.5536$±$\pm 0.02$ $144$
MCTS $(N=4,S=8)$ 0.75750.7575{0.7575}0.7575±$\pm 0.02$ 0.53950.5395{0.5395}0.5395±$\pm 0.04$ 0.96660.9666{0.9666}0.9666±$\pm 0.03$ $3888$
Zero-Order Search $(N=4,K=20)$ $0.6083$±$\pm 0.01$ $0.3751$±$\pm 0.02$ $0.6261$±$\pm 0.04$ $1440$
$\bm{\epsilon}$-greedy ${(N=4,K=20)}$ $\mathbf{0.9813}$±$\pm 0.01$ $\mathbf{0.7208}$±$\pm 0.03$ $\mathbf{0.9885}$±$\pm 0.04$ $1440$
@@
MCTS solves for these drawbacks by approximating exhaustive search over the $N^{T}$ possible root-to-leaf paths in the tree, with the approximation error going to 00 as the number of simulations $S\to\infty$. We indeed see this in Table 1: rejection sampling and beam search beat naive sampling across all reward functions; but excluding zero-order search, MCTS is best across the board, as expected.
@@
Additional experiments, as well as qualitative results, are included in Appendices D and E. Notably, we conduct hyperparameter sweeps and measure how each sampling method scales with increasing NFEs in Appendix D, finding that $\epsilon$-greedy achieves the best scaling laws empirically amongst all sampling methods.
@@ 5.2 Extending to Text-to-Image Generation
Table 2: Stable Diffusion results by sampling method. Following Ma et al. (2025), we use natural language versions of the same class labels used in EDM generation (e.g. “a toucan”). Generating each sample took $\sim$1111 second for naive sampling (the lowest-compute method) and $<$2 minutes for MCTS (the highest-compute method) on a single A100 (40GB). We again provide as error bars the standard deviation of scores for a given prompt over $20$ generations with variability due to randomness of the noise draws.
@@
Naive Sampling $0.4800$±$\pm 0.02$ $0.6314$±$\pm 0.03$ $0.2457$±$\pm 0.00$ $18$
Best of $4$ Sampling $0.5160$±$\pm 0.04$ $0.6847$±$\pm 0.02$ $0.2673$±$\pm 0.00$ $72$
Beam Search ($N=4,B=2$) $0.4981$±$\pm 0.01$ $0.7014$±$\pm 0.01$ $0.2682$±$\pm 0.00$ $144$
MCTS ($N=4,S=8$) 0.60770.6077{0.6077}0.6077±$\pm 0.03$ 0.73220.7322{0.7322}0.7322±$\pm 0.02$ $0.2885$±$\pm 0.01$ 38883888{3888}3888
Zero-Order Search ($N=4,K=20$) $0.5764$±$\pm 0.02$ $0.7281$±$\pm 0.06$ $0.2689$±$\pm 0.02$ $1440$
$\bm{\epsilon}$-greedy ($N=4,K=20$) $\mathbf{0.7110}$±$\pm 0.03$ $\mathbf{0.8052}$±$\pm 0.05$ $\mathbf{0.2973}$±$\pm 0.02$ $1440$
@@ 6 Conclusion
We present the first practical framework for test-time noise trajectory optimization in SDE-based diffusion models. We first cast denoising as an MDP, enabling us to apply sophisticated step-level search algorithms including MCTS; then, via relaxing the MDP to a sequence of independent contextual bandits, we design a novel $\epsilon$-greedy search algorithm that adapts the search strategy (local exploitation vs. global exploration) based on the diffusion step, significantly outperforming vanilla sampling and even MCTS. We extensively discuss broader impacts, limitations, and future work in Appendices F and G.
@@ Appendix C Search Algorithms
1:Discretization timesteps $T$, context (e.g. class label) $\mathbf{c}$, learned denoising network $D_{\theta}$, branching factor $N$, sampling step function $f$, $[0,1]$-bounded reward function $r$, fixed per-step noise candidates $\{\mathbf{z}_{t}^{(j)}\}_{(t,j)\in\{1,\dots,T\}\times\{1,\dots,N\}}$, number of simulations $S$, exploration constant $C=1.414$, initial noise sample $\mathbf{x}_{T}$ Define UCB score($\mathbf{x}_{t},t,\mathbf{z}_{t}^{(j)}$) = reward from ($\frac{\text{reward from }(f(\mathbf{x}_{t},t,\mathbf{c},\mathbf{z}_{t}^{(j)}),t-1)}{\text{\# visits to }(f(\mathbf{x}_{t},t,\mathbf{c},\mathbf{z}_{t}^{(j)}),t-1)}+C\sqrt{\frac{\log\left(\text{\# visits to }(\mathbf{x}_{t},t)\right)}{\text{\# visits to }(f(\mathbf{x}_{t},t,\mathbf{c},\mathbf{z}_{t}^{(j)}),t-1)}}$ $\mathcal{C}(\mathbf{x}_{t})\triangleq$ set of child nodes of $\mathbf{x}_{t}$, s.t. $|\mathcal{C}(\mathbf{x}_{t})|\in\{0,N\}$, $c_{j}(\mathbf{x}_{t})\triangleq f(\mathbf{x}_{t},t,\mathbf{c},\mathbf{z}_{t}^{(j)})$
2:for $t=T$ to $1$ do
@@
8:         for $t^{\prime\prime}=t^{\prime}-1$ to $T$ do                                                         // backpropagation        visits to $(\mathbf{x}_{t^{\prime\prime}},t^{\prime\prime})$ += $1$        reward from $(\mathbf{x}_{t^{\prime\prime}},t^{\prime\prime})$ += $r(\hat{\mathbf{x}}_{0},\mathbf{c})$
@@
2:for $n=1$ to $N$ do
3:     for $t=T$ to $1$ do     Sample noise candidate $\mathbf{z}_{t}^{(n)}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$
@@
2:for $t=T$ to $1$ do Sample $N$ noise candidates $\{\mathbf{z}_{t}^{(i)}\}_{i=1}^{N}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ $\mathcal{C}=\emptyset$
@@
4:         for $i=1$ to $N$ do        Add $f\left(\mathbf{x}_{t}^{(j)},t,\mathbf{c},\mathbf{z}_{t}^{(i)}\right)$ to $\mathcal{C}$
@@
3:for $t=T$ to $1$ do Sample $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$
4:     for $k=1$ to $K$ do Sample $N$ noise candidates $\{\mathbf{z}^{(i)}_{t}\}_{i=1}^{N}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$, for each $i$ let $\mathbf{z}^{(i)}_{t}=\lambda\sqrt{2CHW}\mathbf{z}^{(i)}_{t}+\mathbf{p}$ Set $\mathbf{p}=\mathbf{z}^{(i)}_{t}$ for $i$ s.t. one-step $\hat{\mathbf{x}}_{0}$ prediction using $\mathbf{z}^{(i)}_{t}$ attains highest score under $r$

### v2

@@ Abstract
The iterative and stochastic nature of diffusion models enables test-time scaling, whereby spending additional compute during denoising generates higher-fidelity samples. Increasing the number of denoising steps is the primary scaling axis, but this yields quickly diminishing returns. Instead, optimizing the noise trajectory—the sequence of injected noise vectors—is promising, as the specific noise realizations critically affect sample quality; but this is challenging due to a high-dimensional search space, complex noise-outcome interactions, and costly trajectory evaluations. We address this by first casting diffusion as a Markov Decision Process (MDP) with a terminal reward, showing tree-search methods such as Monte Carlo tree search (MCTS) to be meaningful but impractical. To balance performance and efficiency, we then resort to a relaxation of MDP, where we view denoising as a sequence of independent contextual bandits. This allows us to introduce an $\epsilon$-greedy search algorithm that globally explores at extreme timesteps and locally exploits during the intermediate steps where de-mixing occurs. Experiments on EDM and Stable Diffusion reveal state-of-the-art scores for class-conditioned/text-to-image generation, exceeding baselines by up to $164\%$ and matching/exceeding MCTS performance. To our knowledge, this is the first practical method for test-time noise trajectory optimization of arbitrary (non-differentiable) rewards.
@@ 1 Introduction
Figure 1: (Left) Implicit denoising tree traversed by search algorithms. (Right) Visualization of local search in noise space to maximize reward at a single timestep t{t}.
@@ 2 Related Work
While test-time scaling has been predominantly explored for language models (Yao et al., 2023; Muennighoff et al., 2025), recent work has begun to investigate test-time scaling behavior for diffusion. One means of investing more compute in diffusion model sampling is increasing the number of denoising steps; this generally leads to better generations, but with diminishing benefits, due to accumulation of approximation and discretization errors (Xu et al., 2023b). In this work, we fix the denoising steps and explore scaling along an orthogonal axis: compute devoted to searching for more favorable noise candidates during the sampling process.
@@
Zhang et al. (2025) propose T-SCEND, which sequentially performs best-of-$N$ random search and MCTS as denoising proceeds; however this paper does not truly implement standard MCTS as (1) at timestep $t$ they estimate reward by computing the predicted clean sample $\hat{\mathbf{x}}_{0}^{(t)}$ via Tweedie’s formula (Efron, 2011) (which estimates the posterior mean $\mathbb{E}[\mathbf{x}_{0}\mid\mathbf{x}_{t},\mathbf{c}]$ from a noisy sample $\mathbf{x}_{t}$, see Appendix B) rather than doing a complete rollout and then backpropagating reward on the true clean sample, making their method still a greedy search; and (2) their method performs best-of-$N$ search for the majority of denoising steps, only doing their “MCTS” for a small number ($\sim$33-55) of denoising steps near the very end of the sampling process. This version of MCTS can no longer be considered a true approximation of exhaustive tree search.
@@ 3.2 From Markov Decision Processes to Contextual Bandits
[…]
@@ 4.2 Monte-Carlo Tree Search (MCTS)
MCTS approximates exhaustive search over the $N^{T}$ possible noise trajectories, with the approximation error going to 0 as the number of simulations $S\to\infty$. Full MCTS though faces two key drawbacks: (1) it is extremely computationally expensive, with its number of function evaluations (NFEs)—i.e. number of forward passes through $D_{\theta}$—scaling quadratically in the number of denoising steps $T$ (see Table 3 for each sampling method’s NFE formula); and (2) MCTS by definition is restricted to a fixed denoising tree, i.e. at each timestep $t$ we cannot explore “noise space” beyond the $N$ initially determined candidates. This is a critical issue, as the initial $N$ candidates are sampled uniformly from our prior (i.e., $\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$)) as opposed to any higher-reward regions of noise space that might exist. In other words, we have no notion of importance sampling; we are allocating the entire computation cost of running MCTS to samples that in general, likely yield lower-than-optimal rewards (Chatterjee and Diaconis, 2017).
@@ 4.3 Noise Search via Contextual Bandits
Recall that the SDE diffusion formulation requires the injected noise samples at each timestep to be independent (and Normally distributed). Thus, if we assume independence of actions across timesteps, and ignore the impact of the action at timestep $t$ on the state at t1t-1, we can reformulate denoising as a sequence of $T$ independent continuous contextual bandit problems; at each timestep $t$, we have the bandit characterized by:
@@ 4.4 $\epsilon$}-greedy
Inspired by the semi-uniform strategies originally proposed to approximately solve bandit problems (Slivkins, 2024), including the original $\epsilon$-greedy bandit algorithm (Sutton and Barto, 1998), we introduce a novel sampling algorithm that is an $\epsilon$-contaminated version of zero-order search (henceforth simply $\epsilon$-greedy” (Alg. 1), where at each timestep $t$: (1) We start with a random Gaussian noise $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ as our pivot. (2) We then find $N$ candidates $\{\mathbf{z}_{t}^{(i)}\}_{i=1}^{N}$, via the following procedure: with probability $\epsilon$, $\mathbf{z}_{t}^{(i)}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$; else, we draw a random $\mathbf{z}_{t}^{(i)}$ from the pivot’s neighborhood ($\lambda\sqrt{2d}$-radius ball around $\mathbf{p}$). (3) We compute the clean sample predicted $\hat{\mathbf{x}}_{0}^{(i)}$ using Tweedie’s formula using each candidate $\mathbf{z}_{t}^{(i)}$. (4) Finally, we score each clean sample, set the pivot to the noise candidate leading to the clean sample with the highest scoring, and repeat steps 1-3 K1K-1 times, finally setting $\mathbf{z}_{t}$ to the last value of $\mathbf{p}$.
@@
1:Discretization timesteps $T$, context (e.g. class label) $\mathbf{c}$, learned denoising network $D_{\theta}$, max. step size scaling factor $\lambda$, number of local search iterations $K$, branching factor (number of noise candidates) NN, sampling step function $f$, mixture proportion $\epsilon$, reward function $r$, initial noise sample $\mathbf{x}_{T}$
2:for $t=T$ to 11 do Sample $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ // pivot
@@
yields remarkable performance improvements, which we discuss in section 5. We also provide a theoretical formalization of our algorithm’s effectiveness in Appendix D.
@@ 5.1.3 Results
Table 1: EDM results by sampling method. Each value in columns 2-4 is obtained by generating $36$ images given random ImageNet class labels. The column header denotes the reward used both to score the final images and during sampling if applicable. We set $\lambda=0.15$ and $\epsilon=0.4$. Note that we measure distance as $\|\mathbf{z}_{t}^{(i)}-\mathbf{p}\|_{F}$, hence let $\lambda$ a scaling factor applied to $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{N}(0,1)^{d}}\left[\|\mathbf{x}-\mathbf{y}\|_{F}\right]\approx\sqrt{2d}$ (where $\|\cdot\|_{F}$ is the Frobenius norm. Re. notation, NN is branching factor, $B$ is beam width, $S$ is number of MCTS simulations, and $K$ is the number of local search iterations. Generating each sample took <<11 second for naive sampling (the lowest-compute method) and <<1 minute for MCTS (the highest-compute method) on a single A100 (40GB). We provide error bars for each reward and sampling method, computed as the standard deviation of scores for a given prompt over 2020 generations with variability from randomness of the noise draws.
@@
[…]
$\bm{\epsilon}$-greedy (N=4,K=20){(N=4,K=20)} $\mathbf{0.9813}$$\pm 0.01$ $\mathbf{0.7208}$$\pm 0.03$ $\mathbf{0.9885}$$\pm 0.04$ $1440$
@@
MCTS solves for these drawbacks by approximating exhaustive search over the $N^{T}$ possible root-to-leaf paths in the tree, with the approximation error going to 0 as the number of simulations $S\to\infty$. We indeed see this in Table 1: rejection sampling and beam search beat naive sampling across all reward functions; but excluding zero-order search, MCTS is best across the board, as expected.
@@
Additional experiments, as well as qualitative results, are included in Appendices F, G and H. Notably, we conduct hyperparameter sweeps and measure how each sampling method scales with increasing NFEs in Appendix G, finding that $\epsilon$-greedy achieves the best scaling laws empirically amongst all sampling methods.
@@ 5.2 Extending to Text-to-Image Generation
Table 2: Stable Diffusion results by sampling method. Following Ma et al. (2025), we use natural language versions of the same class labels used in EDM generation (e.g. “a toucan”). Generating each sample took $\sim$11 second for naive sampling (the lowest-compute method) and <<2 minutes for MCTS (the highest-compute method) on a single A100 (40GB). We again provide as error bars the standard deviation of scores for a given prompt over 2020 generations with variability due to randomness of the noise draws.
@@
[…]
$\bm{\epsilon}$-greedy ($N=4,K=20$) $\mathbf{0.7110}$$\pm 0.03$ $\mathbf{0.8052}$$\pm 0.05$ $\mathbf{0.2973}$$\pm 0.02$ $1440$
@@ 6 Conclusion
We present the first practical framework for test-time noise trajectory optimization in SDE-based diffusion models. We first cast denoising as an MDP, enabling us to apply sophisticated step-level search algorithms including MCTS; then, via relaxing the MDP to a sequence of independent contextual bandits, we design a novel $\epsilon$-greedy search algorithm that adapts the search strategy (local exploitation vs. global exploration) based on the diffusion step, significantly outperforming vanilla sampling and even MCTS. We extensively discuss broader impacts, limitations, and future work in Appendices I and J.
@@ Appendix C Search Algorithms
[…]
2:for $t=T$ to 11 do
@@
8:   for $t^{\prime\prime}=t^{\prime}-1$ to $T$ do                       // backpropagation    visits to $(\mathbf{x}_{t^{\prime\prime}},t^{\prime\prime})$ += 11    reward from $(\mathbf{x}_{t^{\prime\prime}},t^{\prime\prime})$ += $r(\hat{\mathbf{x}}_{0},\mathbf{c})$
@@
2:for $n=1$ to NN do
3:  for $t=T$ to 11 do     Sample noise candidate $\mathbf{z}_{t}^{(n)}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$
@@
2:for $t=T$ to 11 do Sample NN noise candidates $\{\mathbf{z}_{t}^{(i)}\}_{i=1}^{N}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$ $\mathcal{C}=\emptyset$
@@
4:   for $i=1$ to NN do    Add $f\left(\mathbf{x}_{t}^{(j)},t,\mathbf{c},\mathbf{z}_{t}^{(i)}\right)$ to $\mathcal{C}$
@@
3:for $t=T$ to 11 do Sample $\mathbf{p}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$
4:  for $k=1$ to $K$ do Sample NN noise candidates $\{\mathbf{z}^{(i)}_{t}\}_{i=1}^{N}\sim\mathcal{N}(\mathbf{0},\sigma_{t}^{2}\mathbf{I})$, for each $i$ let $\mathbf{z}^{(i)}_{t}=\lambda\sqrt{2CHW}\mathbf{z}^{(i)}_{t}+\mathbf{p}$ Set $\mathbf{p}=\mathbf{z}^{(i)}_{t}$ for $i$ s.t. one-step $\hat{\mathbf{x}}_{0}$ prediction using $\mathbf{z}^{(i)}_{t}$ attains highest score under $r$